### v2

@@ 1 Introduction
$\max_{t\leq\overline{t}}\ E[\text{benefit}\cdot\mathbf{1}\{\text{income}\leq t\}]$
@@
and approximates the optimal threshold based on benefits estimated from experimental data. Recent papers have argued the EWM approach is attractive under a wide range of data distributions, including Athey and Wager (2021), among others. Notably the average benefits under the eligibility policy selected by EWM converges to the highest attainable level as the sample size grows. This property can be thought of as uniform asymptotic welfare efficiency.
@@
In practice policymakers often face budget constraints, but only have imperfect information about whether a given eligibility policy satisfies the budget constraint. First, there may be imperfect take-up: eligible individuals might not participate in the welfare program, resulting in zero cost to the government, e.g. Finkelstein and Notowidigdo (2019). Second, costs incurred by eligible individuals who participate in the welfare program may vary considerably, largely driven by individuals’ different needs but also many other factors, e.g. Finkelstein et al. (2017). Both considerations are hard to predict ex ante, implying that the potential cost of providing eligibility to any given individual is unknown at the time of designing the eligibility policy. Unobservability of the potential cost requires estimation based on experimental data, contributing to uncertainty in the budget estimate of a given eligibility policy.
For this empirically relevant setting where the budget needed to implement an eligibility policy involves an unknown cost, this paper introduces a new property of statistical rules, namely asymptotic feasibility. Policies are feasible if they satisfy the budget constraint in the target population; otherwise, they are referred to as infeasible. A statistical rule is asymptotically feasible if given a large enough experimental sample, the statistical rule is very likely to select feasible eligibility policies. Since a potential cost overrun due to an underestimate of the budget can be very costly to policymakers, asymptotic feasibility is also a desirable property in the current setting, especially for policymakers who are extremely cautious about cost overruns. This paper answers three questions in the current setting with unknown cost: is there any statistical rule that achieves uniformly good performance for a wide range of data distributions as in the original EWM literature, whether the obvious extension of the existing EWM statistical rule remains attractive, and what are some alternative statistical rules.
@@
Second, I show the direct extension of the existing EWM statistical rule is not appealing in the setting with unknown cost. The reason is that this sample-analog rule ignores the estimation error in the estimated budget constraint, which has non-negligible consequences even when the sample size is large. For data distributions mentioned earlier where the budget constraint is exactly binding, the welfare loss does not vanish with sample size. The probability of selecting infeasible policies also does not vanish with sample size, but a simple modification to the sample-analog rule can ensure that feasible policies are selected with high probability. These considerations are relevant in real-world settings, as demonstrated by a calibrated simulation study.
Finally I propose the trade-off rule, which can select infeasible eligibility policies, but only if borrowing money and exceeding the budget constraint are not too costly to justify the marginal gain in welfare from violating the budget constraint. I show the trade-off rule is uniformly asymptotically welfare efficient, which is an improvement over the direct extension of the EWM rule.
@@
To illustrate the statistical rules proposed in this paper, I consider a budget-constrained Medicaid expansion. Medicaid is a government-sponsored health insurance program intended for the low-income population in the United States. Based on the Oregon Health Insurance Experiment (OHIE) conducted in 2008, I derive Medicaid expansion policies using statistical rules. Oregon’s current Medicaid expansion policy is based on a threshold for household income only. In my application, I examine whether health can be improved by allowing the income threshold to vary by the number of children in the household, setting the budget constraint to the current level. Using the sample-analog rule, the selected policy limits eligibility to be lower than the current level to meet the estimated budget constraint. However, the cost estimates are highly variable, indicating considerable uncertainty about whether the selected policy is welfare-efficient in the population. In contrast, the trade-off rule selects an eligibility policy that expands eligibility for many households above the current level, especially those with children. The higher level occurs because, based on the OHIE data, Medicaid improves health for these households, and the additional health benefit from violating the budget constraint exceeds the cost of doing so, assuming a reasonable upper bound on the monetary value on the health benefit.
@@
The rest of the paper proceeds as follows. Section 1.1 discusses related work in more detail. Section 2 presents theoretical results. Sections 3 and 4 discuss properties of two statistical rules. Section 5 conducts a simulation study and Section 6 considers an empirical example of designing a more flexible Medicaid expansion policy for the low-income population in Oregon. Section 7 concludes. Proofs can be found in the Appendix. Supporting lemmas, additional results and computational details can be found in the Online Appendix.
@@ 1.1 Literature review
The literature on statistical rules in econometrics has also developed a definition for optimality of statistical rules. Manski (2004) considers the minimax-regret, defined to be loss in expected welfare achieved by the statistical rule relative to the welfare achieved by the theoretically optimal eligibility policy. In the absence of any constraint, under the theoretically optimal eligibility policy, anyone with positive benefit from the welfare program would be assigned with eligibility. The minimax regret is the upper bound on the regret that results from not knowing the data distribution. Manski (2004) argues a statistical rule is preferred if its minimax-regret converges to zero with the sample size, and analyzes the minimax-regret for the class of statistical rules that only selects eligibility policy based on subsets of the observed covariates. Stoye (2009) shows that with continuous covariates and no functional form restrictions on the set of policies, minimax regret does not converge to zero with the sample size because the theoretically optimal policy can be too difficult to approximate by statistical rule. Kitagawa and Tetenov (2018) avoid this issue by imposing functional form restrictions. They propose the EWM rule, which starts with functional form restrictions on the class of available policies, and then selects the policy with the highest estimated benefit (empirical welfare) based on an RCT sample. They prove the optimality of EWM in the sense that its regret converges to zero at the minimax rate. Importantly, the regret is defined to be loss in expected welfare relative to the maximum achievable welfare in the constrained class, which avoids the negative results of Stoye (2009). Athey and Wager (2021) propose doubly-robust estimation of the average benefit, which leads to an optimal rule even with quasi-experimental data. Mbakop and Tabord-Meehan (2021) propose a Penalized Welfare Maximization assignment rule which relaxes restrictions of the policy class.
@@ 2.1 Motivation and setup
I begin by setting up a general constrained optimization problem, which depends on the following random attributes of an individual:
$A=(\Gamma,R,X)\in\mathcal{A}\subseteq\mathbb{R}^{2+p}.$ (1)
Here $\Gamma$ is benefit from the treatment for the individual, $R$ is the cost to the policymaker of providing the individual with the treatment, and $X\in\mathcal{X}\subset\mathbb{R}^{p}$ denotes their $p$-dimensional characteristics. The individual belongs to a population that can be characterized by the joint distribution $P$ on the random attributes $A$. The unknown distribution $P\in\mathcal{P}$ is from a class of distributions $\mathcal{P}$.
@@
$\max_{g\in\mathcal{G}}E_{P}[\Gamma\cdot g(X)]\ s.t.\ E_{P}[R\cdot g(X)]$ $\leq k.$ (2)
If the policymaker does not have a fixed budget but still want to account for cost, the scalar $\Gamma$ can be the difference in benefit and cost. Since a fixed budget is common, I impose a harsh constraint.
The benefit-cost attributes $(\Gamma,R)$ of any given individual may or may not be observed. When the benefit-cost attributes are observed, policymakers observe i.i.d. $A_{i}=(\Gamma_{i},R_{i},X_{i})$ of individual $i$ in a random sample of the target population. When the benefit-cost attributes are unobserved, the focus of this paper is the setting where policymakers can construct their estimates $(\Gamma_{i}^{\ast},R_{i}^{\ast})$ in a random sample of sample size $n$ along with the characteristics $X_{i}$ from an experiment or quasi-experiment that satisfy Assumption 3.1, as discussed later.
@@
$\max_{g\in\mathcal{G}}E_{P}[E_{P}[\Gamma\mid X]\cdot g(X)]\ s.t.\ E_{P}[E_{P}[R\mid X]\cdot g(X)]\leq k.$ (3)
When the eligibility policy can be based on any characteristics whatsoever, the class of available policies is unrestricted i.e. $\mathcal{G}=2^{\mathcal{X}}$. In this unrestricted class, when the cost is non-negative, the above expression makes clear that the optimal eligibility policy is based on thresholding by the benefit-cost ratio $E_{P}[\Gamma\mid X]/E_{P}[R\mid X]$ where the numerator and the denominator are respectively the average effects conditional on the observed characteristics (CATE) and the conditional average resource required. Online Appendix B provides a formal statement.
@@
I next specialize the constrained optimization problem to selecting eligibility policy for welfare programs with the example of Medicaid expansion. In the example of implementing welfare programs, policies take the form of eligibility policies. I restrict attention to non-randomized policies as in the leading example of welfare programs, deterministic policies such as income thresholds are more relevant. While the optimization problem policymakers face in practice can take on different forms, constrained optimization problems of this mathematical structural are ubiquitous in social science. Theoretically oriented readers may proceed directly to Section 2.2.
@@
Suppose the government wants to implement some welfare program. The treatment in this example is eligibility for such welfare program. Due to a limited budget, the government cannot make eligibility universal and can only provide eligibility to a subpopulation. To use the budget efficiently, policymakers consider the constrained optimization problem (2). In this example, the policy $g(X)$ assigns an individual to eligibility based on their observed characteristics $X$, and is usually referred to as an eligibility policy. I denote $\Gamma$ to be the benefit experienced by an individual after receiving eligibility for the welfare program. Specifically, let $(Y_{1},Y_{0})$ denote the potential outcomes that would have been observed if an individual were assigned with and without eligibility, respectively. The benefit from eligibility policy is therefore defined as $\Gamma\coloneqq Y_{1}-Y_{0}$. Note that maximizing benefit is equivalent to maximizing the outcomes (welfare) under the utilitarian social welfare function: $E_{P}[Y_{1}\cdot g(X)+Y_{0}(1-g(X))]$. I denote $R$ to be the potential cost from providing an individual with eligibility for the welfare program. Both $\Gamma$ and $R$ are unobserved at the time of assignment and will need to be estimated.
@@
Suppose policymakers want to maximize the health benefit of Medicaid by adopting a more flexible expansion policy. Specifically, they relax the uniform income threshold of 138% and allows the income thresholds to vary with the number of children in the household. It may be infeasible to implement a new expansion policy if it costs more than the current one. Therefore I impose the constraint that the more flexible policy must cost no more than the current one.
Correspondingly, the constrained optimization problem (2) sets $\Gamma$ to be the health benefit from Medicaid, $R$ to be the excess per-enrollee cost of Medicaid relative to the current expansion policy, and the appropriate threshold to be $k=0$. The characteristics $X$ include both income and number of children in the household.
@@ 2.2 Desirable properties for statistical rules
$W(g;P)=E_{P}[\Gamma\cdot g(X)],\ B(g;P)=E_{P}[R\cdot g(X)].$ (5)
@@
When the benefit-cost attributes $(\Gamma,R)$ are unobserved and the distribution $P$ is unknown, both the welfare function and the budget function are unknown functions. Denote by $\widehat{g}$ a statistical rule that selects an eligibility policy after observing some experimental data of sample size $n$ distributed according to $P^{n}$. This section provides formal definitions for two desirable properties of $\widehat{g}$.
@@
$\lim_{n\rightarrow\infty}Pr_{P^{n}}\left\{W(\widehat{g};P)-W(g_{P}^{\ast};P)<-\epsilon\right\}=0,$ (6)
@@
$\lim_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}Pr_{P^{n}}\left\{W(\widehat{g};P)-W(g_{P}^{\ast};P)<-\epsilon\right\}=0.$
@@
$\lim_{n\rightarrow\infty}Pr_{P^{n}}\{B(\widehat{g};P)>k\}=0,$
@@
$\lim_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}Pr_{P^{n}}\{B(\widehat{g};P)>k\}=0.$ (7)

### v3

@@ 1 Introduction
$\max_{t\leq\overline{t}}\ \mathbb{E}[\text{benefit}\cdot\mathbf{1}\{\text{income}\leq t\}]$
@@
and approximates the optimal threshold based on benefits estimated from experimental data. Recent work has demonstrated that the EWM approach performs well across a broad range of data distributions. As the sample size grows, the average of benefits obtained under the eligibility policy selected by EWM converges to the highest attainable level, a property I refer to as as uniform asymptotic welfare efficiency.
@@
I follow this line of work in focusing on settings where the eligibility policy for welfare programs must be determined ex ante and cannot be easily adjusted during implementation, such as Food Stamp and Medicaid in the United States. [^1] There are also settings in which eligibility is implemented sequentially until the budget is exhausted, such as anti-poverty programs in Chile as analyzed by Carneiro et al. (2019). My analysis does not extend to those cases, as it does not account for the possibility that the benefits and costs may vary depending on the order in which individuals enroll. In practice policymakers often face budget constraints, but only have imperfect information about whether a given eligibility policy satisfies the budget constraint. First, there may be imperfect take-up: eligible individuals might not participate in the welfare program, resulting in zero cost to the government, e.g. Finkelstein and Notowidigdo (2019). Second, costs incurred by eligible individuals who participate in the welfare program may vary considerably, largely driven by individuals’ different needs but also many other factors, e.g. Finkelstein et al. (2017). Both considerations are hard to predict ex ante, implying that the potential cost of providing eligibility to any given individual is unknown at the time of designing the eligibility policy. Unobservability of the potential cost requires estimation based on experimental data, contributing to uncertainty in the budget estimate of a given eligibility policy.
For this empirically relevant setting where the budget needed to implement an eligibility policy involves an unknown cost, this paper introduces a new property of statistical rules, namely asymptotic feasibility. Policies are feasible if they satisfy the budget constraint in the target population; otherwise, they are referred to as infeasible. A statistical rule is asymptotically feasible if given a large enough experimental sample, the statistical rule is very likely to select feasible eligibility policies. While budget overruns are generally tolerated in countercyclical programs like unemployment insurance and Medicaid in the United States, in other settings, such as subsidized health and education programs in developing countries, a potential budget overrun due to an underestimated budget can be highly undesirable for policymakers, as securing additional funding may be difficult. [^2] For example, international and national funding for malaria control has fallen short of what is estimated to be needed in recent years. Motivated by the fact that Kenya can afford to distribute bed net subsidies to only 50% of its target population in 2007, Bhattacharya and Dupas (2012) analyzed the constrained optimal allocation to increase take-up. In this context, ensuring asymptotic feasibility is particularly desirable, especially for policymakers who are highly risk-averse to budget overruns.
This paper answers three questions in the current setting with unknown cost: is there any statistical rule that achieves uniformly good performance for a wide range of data distributions in terms of both welfare efficiency and feasibility, whether the obvious extension of the existing EWM statistical rule remains attractive, and what are some alternative statistical rules.
@@
Second, I show the direct extension of the existing EWM statistical rule is not appealing in the setting with unknown cost. The reason is that this sample-analog rule ignores the estimation error in the estimated budget needed to implement a given policy, which has non-negligible consequences even when the sample size is large. For data distributions mentioned earlier where the budget constraint is exactly binding, the welfare loss does not vanish with sample size. The probability of selecting infeasible policies also does not vanish with sample size. Intuitively, this latter issue can be mitigated by using a slightly downward-biased version of the budget constraint, effectively making the budget estimate more conservative. [^3] Similar intuition arises in optimal prediction under asymmetric loss (Christoffersen and Diebold, 1997), where positive prediction errors are more costly than negative ones, making a downward-biased predictor optimal. I formalize this idea by setting the degree of conservativeness proportional to the standard error of the budget estimate, effectively selecting only policies with an upper confidence bound below the budget constraint. This modification to the sample-analog rule ensures that feasible policies are selected with high probability.
So far, the discussion assumes that even minor budget violations are unacceptable. However, in practice, exceeding the budget constraint may be desirable if the welfare gains outweigh the borrowing costs. [^4] What happens when a welfare program goes over budget is highly context-specific, and there are various ways to model it. In this paper, I focus specifically on the scenario where the government borrows money to cover the excess cost while other models-such as those involving rationing as in Kitagawa and Tetenov (2018) and poorer services are certainly possible. To account for this situation, I introduce a new objective function that maximizes welfare gains but imposes some penalty once spending exceeds the budget, thereby giving the new objective function the intuitive interpretation of a trade-off function. I propose the trade-off rule, which optimizes the sample-analog version of the trade-off function. I show the trade-off rule is uniformly asymptotically welfare efficient.
@@
To illustrate the trade-off rule, I apply it to data from the Oregon Health Insurance Experiment (OHIE) to select a more flexible Medicaid expansion policy than the current one. Medicaid is a government-sponsored health insurance program intended for the low-income population in the United States. The current Medicaid expansion policy determines eligibility solely based on household income. I examine whether health can be improved by allowing the income threshold to vary by the number of children in the household, setting the budget constraint equal to the cost of the current policy. Imposing a reasonable penalty for exceeding the budget in this context, that any overrun needs to be repaid in full, the trade-off rule selects an eligibility policy that expands eligibility for many households above the current threshold, especially those with children. This occurs because, based on the OHIE data, the additional health benefit from extending eligibility to these households outweigh the penalty of having to repay the overrun in full.
@@
The rest of the paper proceeds as follows. Section 1.1 discusses related work in more detail. Section 2 presents theoretical results. Sections 3 and 4 discuss properties of two statistical rules, illustrated by an empirical example of designing a more flexible Medicaid expansion policy for the low-income population in Oregon. Section 5 conducts a simulation study and Section 6 concludes. Proofs, supporting lemmas, additional results and computational details can be found in the Appendix.
@@ 1.1 Literature review
The literature on statistical rules in econometrics has also developed a definition for optimality of statistical rules. Manski (2004) considers the regret, defined to be loss in expected welfare achieved by the statistical rule relative to the welfare achieved by the theoretically optimal eligibility policy. In the absence of any constraint, under the theoretically optimal eligibility policy, anyone with positive benefit from the welfare program would be assigned with eligibility. The minimax regret rule minimizes the upper bound on the regret that results from not knowing the data distribution. Stoye (2009) shows that with continuous covariates and no functional form restrictions on the set of policies, minimax regret does not converge to zero with the sample size because the theoretically optimal policy can be too difficult to approximate by a statistical rule. Kitagawa and Tetenov (2018) avoid this issue by imposing functional form restrictions. They propose the EWM rule, which starts with functional form restrictions on the class of available policies, and then selects the policy with the highest estimated benefit (empirical welfare) based on an RCT sample. They prove the optimality of EWM in the sense that its regret converges to zero at the minimax rate. Importantly, the regret is defined to be loss in expected welfare relative to the maximum achievable welfare in the constrained class, which avoids the negative results of Stoye (2009). Athey and Wager (2021) propose doubly-robust estimation of the average benefit, which leads to an optimal rule even with quasi-experimental data. Mbakop and Tabord-Meehan (2021) propose a Penalized Welfare Maximization rule which relaxes restrictions of the policy class.
@@ 2.1 Motivation and setup
I begin by setting up a general constrained optimization problem, which depends on the following attributes of an individual:
$A=(\tau,C,X)\in\mathcal{A}\subseteq\mathbb{R}^{2+p}.$
Here $\tau$ is benefit from the treatment for the individual, $C$ is the cost to the policymaker of providing the individual with the treatment, and $X\in\mathcal{X}\subset\mathbb{R}^{p}$ denotes their $p$-dimensional characteristics. The individual belongs to a population that can be characterized by the joint distribution $P$ on the attributes $A$. The unknown distribution $P\in\mathcal{P}$ is from a class of distributions $\mathcal{P}$.
@@
$\max_{g\in\mathcal{G}}\mathbb{E}_{P}[\tau\cdot g(X)]\text{ subject to }\mathbb{E}_{P}[C\cdot g(X)]$ $\leq k.$ (1)
If the policymaker does not have a fixed budget but still wants to account for cost, the scalar $\tau$ can be the difference in benefit and cost. I impose a harsh constraint at known $k$ to model a fixed budget.
The benefit-cost attributes $(\tau,C)$ of any given individual may be unobserved in practice. The focus of this paper is the setting where policymakers can construct their estimates $(\tau_{i}^{\ast},C_{i}^{\ast})$ in a random sample of sample size $n$ along with the characteristics $X_{i}$ from an experiment or quasi-experiment that satisfy Assumption 3.1, as discussed later.
@@
$\max_{g\in\mathcal{G}}\mathbb{E}_{P}[\mathbb{E}_{P}[\tau\mid X]\cdot g(X)]\text{ subject to }\mathbb{E}_{P}[\mathbb{E}_{P}[C\mid X]\cdot g(X)]\leq k.$
When the eligibility policy can be based on any characteristics whatsoever, the class of available policies is unrestricted i.e. $\mathcal{G}=2^{\mathcal{X}}$. In this unrestricted class, when the cost is non-negative, the above expression makes clear that the optimal eligibility policy is based on thresholding by the benefit-cost ratio $\mathbb{E}_{P}[\tau\mid X]/\mathbb{E}_{P}[C\mid X]$ where the numerator and the denominator are respectively the average effects conditional on the observed characteristics (CATE) and the conditional average resource required. Appendix B provides a formal statement.
@@
I next specialize the constrained optimization problem to selecting eligibility policy for welfare programs with the example of Medicaid expansion. In the example of implementing welfare programs, policies take the form of eligibility policies. I restrict attention to non-randomized policies as in the leading example of welfare programs, deterministic policies such as income thresholds are more relevant. Theoretically oriented readers may proceed directly to Section 2.2.
@@
Suppose the government wants to implement some welfare program. The treatment in this example is eligibility for such welfare program. Due to a limited budget, the government cannot make eligibility universal and can only provide eligibility to a subpopulation. To use the budget efficiently, policymakers consider the constrained optimization problem (1). In this example, the policy $g(X)$ assigns an individual to eligibility based on their observed characteristics $X$, and is usually referred to as an eligibility policy. I denote $\tau$ to be the benefit experienced by an individual after receiving eligibility for the welfare program. Specifically, let $(Y_{1},Y_{0})$ denote the potential outcomes that would have been observed if an individual were assigned with and without eligibility, respectively. The benefit from eligibility policy is therefore defined as $\tau\coloneqq Y_{1}-Y_{0}$. Note that maximizing benefit is equivalent to maximizing the outcomes (welfare) under the utilitarian social welfare function: $\mathbb{E}_{P}[Y_{1}\cdot g(X)+Y_{0}(1-g(X))]$. I denote $C$ to be the potential cost from providing an individual with eligibility for the welfare program. Both $\tau$ and $C$ are unobserved at the time of assignment and will need to be estimated.
@@
Suppose policymakers want to maximize the health benefit of Medicaid by adopting a more flexible expansion policy. Specifically, they relax the uniform income threshold of 138% and allow the income thresholds to vary with the number of children in the household. Once the income thresholds are set, they must be codified in state legislation and publicly announced. [^6] The legislated eligibility policy is publicly available on federal websites such as MACPAC. Therefore, in this example, it is reasonable to assume that the eligibility policy must be determined ex ante.
@@
Convincing policymakers to adopt a more flexible expansion policy as in (2) may still require specifying a clear budget target to ensure that the new policy does not exceed the expenditure level of the current expansion with the uniform income threshold of 138%. Given that Medicaid is a countercyclical program and has some flexibility to accommodate budget overruns, later in Section 4, I develop a new rule to allow explicit trade-off between welfare gains from additional spending and a penalty for budget overruns.
Correspondingly, the constrained optimization problem (1) sets $\tau$ to be the health benefit from Medicaid, $C$ to be the cost to Medicaid, and the appropriate threshold $k$ to be the average cost to Medicaid under the current expansion policy with the uniform income threshold of 138%. The characteristics $X$ include both income and number of children in the household.
@@ 2.2 Desirable properties for statistical rules
$W(g;P)=\mathbb{E}_{P}[\tau\cdot g(X)],\ B(g;P)=\mathbb{E}_{P}[C\cdot g(X)].$
@@
When the benefit-cost attributes $(\tau,C)$ are unobserved and the distribution $P$ is unknown, both the welfare function and the budget function are unknown functions. Denote by $\widehat{g}$ a statistical rule that selects an eligibility policy after observing some experimental data of sample size $n$ distributed according to $P^{n}$. This section provides formal definitions for two desirable properties of $\widehat{g}$.
@@
$\lim_{n\rightarrow\infty}\operatorname{Pr}_{P^{n}}\left\{W(\widehat{g};P)-W(g_{P}^{\ast};P)<-\epsilon\right\}=0,$
@@
$\lim_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}\operatorname{Pr}_{P^{n}}\left\{W(\widehat{g};P)-W(g_{P}^{\ast};P)<-\epsilon\right\}=0.$
@@
$\lim_{n\rightarrow\infty}\operatorname{Pr}_{P^{n}}\{B(\widehat{g};P)>k\}=0,$
@@
$\lim_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}\operatorname{Pr}_{P^{n}}\{B(\widehat{g};P)>k\}=0.$